OpenStudy (anonymous):

Jeff spent 1/2 on his Saturday earnings on a pair of shoes and 1/2 of the remaining money on a CD. After he spent $5.35 on lunch, he had $10.85 left. How much did he earn Saturday? A.$32.25 B.$36.45 C.$60.40 D.$64.80 Please explain.

OpenStudy (lgbasallote):

reverse all the operations..change subtraction to addition..division to multiplication ((10.85 + 5.35)2)2 =? i'll leave that to you :P

OpenStudy (anonymous):

Thanks, wait please, I will type my answer shortly.

OpenStudy (lgbasallote):

okay

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Order of operations, PEMDAS? (Parentheses Exponents Multiplying Dividing Adding Subtracting)?

OpenStudy (anonymous):

Do I use this?

OpenStudy (lgbasallote):

right now yes :D

OpenStudy (anonymous):

ok. Wait 1 sec.

OpenStudy (anonymous):

$48.75

OpenStudy (lgbasallote):

hmm seems wrong...try again :) and also that wasnt in the choices haha

OpenStudy (anonymous):

(embarrased) lol

OpenStudy (anonymous):

Please click on the image,it show the online calculator i was using.

OpenStudy (lgbasallote):

4 x 10.85 + 5.35 is not equal to what i wrote...according to this you'll multiply 4 by 10.85 first...but you need to add first THEN multiply..remember i had a parenthesis

OpenStudy (anonymous):

Were did you get the 4 from?

OpenStudy (anonymous):

Btw, I am making a Google Docsl with the problems on my worksheet, is it ok if i give you the online link to it after I finish it?

OpenStudy (anonymous):

lgbaallote?

OpenStudy (anonymous):

basallote (sorry,mispelled your username)

OpenStudy (anonymous):

One way is to use algebra. The other way is to use logic ! First the logical way. Since Jeffrey has spent one half of the amount on a pair of shoes , and one half of the remaining amount on a CD , it means he has spent 3/4 of his earnings. He is therefore left with just a quarter of his earnings. After spending $ 5.35 from this on lunch , he has $ 10.85 left. This means that one quarter of his earnings are $ 10.85 + $ 5.35 = $ 16.20 Thus his total earnings are $ 16.20 * 4 = $ 64.80 In case you use algebra , assume that his earnings are 'x'. His total expenditure is x/2 + 1/2(x - x/2) + 5.35 What he has left with him will be his earnings less his expenditure i.e. x - (x/2 + 1/2(x - x/2) + 5.35). According to the problem , this is 10.85 Equating the two : x - x/2 -x/2 + x/4 - 5.35 = 10.85 i.e. x/4 = 10.85 + 5.35 = 16.20 which gives x = 4 * 16.20 = 64.80.

OpenStudy (lgbasallote):

sorry i went out...^that's the answer

OpenStudy (anonymous):

kumar, I have not learned how to work out those types of equations; could you simplify or shorten your explanation please?

OpenStudy (lgbasallote):

add 10.85 and 5.35 first...it will become 16.20. multiplied by 2 = 32.4..multiply by 2 again...64.8

OpenStudy (lgbasallote):

((10.85 = 5.35)2)2 = ((16.20)2)2 = (32.40)2 = 64.80

OpenStudy (lgbasallote):

^that's + in the parenthesis

OpenStudy (anonymous):

Ok. I get it now, thanks, lg! so ccan I send you the link? it's only a document with my worksheet's problems typed on it.

OpenStudy (lgbasallote):

okay

OpenStudy (anonymous):

tnx.

OpenStudy (anonymous):

im typing as fast as i can.